OpenStudy (anonymous):

A projectile is fired upward from ground level with an initial velocity of 320 per second. (Assume t=0seconds corresponds to the time the object is fired. Use 32ft/sec^2 as acceleration due to gravity.) (s=-32t^2+vOt+sO) At what instant will it be back at ground level? When will the height exceed 1584 feet?

OpenStudy (xishem):

Do you know what s represents in the equation?

OpenStudy (anonymous):

S= hieght of object

OpenStudy (xishem):

Alright, so ground level is 0 height. If you plug in 0 for the height and solve for time, you will find the instant at which the projectile is back at ground level. Keep in mind that you'll get two answers since it is at ground level when you fired it (t=0) and at another time later on.

OpenStudy (anonymous):

0=32t^2+320+0?

OpenStudy (xishem):

Close, just a few mistakes. Be careful about signs and variables:\[0=-32t^2+320t+0\]

OpenStudy (anonymous):

so how do you figuire out the time its back at ground level?

OpenStudy (xishem):

You've put in the conditions, you just need to find the values of t that satisfy the equation.

OpenStudy (anonymous):

-32t(t+10) t=-10 t=0

OpenStudy (anonymous):

sorry negative to t=10

OpenStudy (anonymous):

1584 <-32t^2+320t 0<-32t^2+320t-1584 Then I just use the quadratic formula to figuire it out right?

OpenStudy (xishem):

That's right. Yep.

OpenStudy (anonymous):

Kay thank you. :) you are the best. <3

OpenStudy (xishem):

Are you having troubles?

OpenStudy (anonymous):

sorry it redid that

OpenStudy (dan815):

umm hey

OpenStudy (dan815):

you got a mistake in your equation

OpenStudy (anonymous):

my labtop is being weird but i got the answer \[(-\sqrt{10t+\frac{ 99 }{ 2 }}, \sqrt{10t +\frac{ 99 }{ 2 }})\]

OpenStudy (dan815):